OpenStudy (anonymous):

Consider the curve y = x − x7

OpenStudy (anonymous):

slop to (,10)

OpenStudy (anonymous):

Find the slope of the tangent line to the curve at the point (1, 0).

OpenStudy (anonymous):

dy/dx = 1 - 7x^6 plug in x-value of point (1,0) dy/dx = 1 - 7 = -6